The market for lemon has 10 potential consumers, each having an individual demand curve (5%)

P = 101 - 10Qi, where P is price in dollars per cup and Qi is the number of cups demanded per week by the ith consumer.

a. Find the market demand curve using algebra

b. What is the quantity demanded by each consumer and in the market as a whole when lemon is priced at P = $1/cup?


1
Expert's answer
2021-08-10T18:02:32-0400

Solution:

a.). Individual demand curve: P = 101 – 10Qi

The market demand curve for lemonade is the horizontal summation of the individual demand curves.

Qi = 101/10 – P/10 = 10.1 – 0.1P

Qi = 10.1 – 0.1P

Since we have 10 consumers, we multiply by 10 to get the market demand curve:

10(10.1 – 0.1P) = 101 – P

The market demand curve for lemonade is: Qi = 101 – P


b.). The quantity demanded by each consumer:

Individual demand: Qi = 10.1 – 0.1P

P = 1

Qi = 10.1 – 0.1(1) = 10.1 – 0.1 = 10

The quantity demanded by each consumer = 10

The quantity demanded in the market as a whole:

Qi = 101 – P

Qi = 101 – 1 = 100

Qi = 100

The quantity demanded in the market as a whole = 100